OpenStudy (anonymous):

In a geometric sequence {an}, if a1 = 4 and r = 0.2, the first 4 terms in the sequence are

jimthompson5910 (jim_thompson5910):

a2 = a1*r a2 = 4*0.2 a2 = ???